OpenStudy (chrisplusian):

Need help understanding how an example from my textbook represented a trigonometric fourier series as a compact trigonometric fourier series..... I will type the example out in the comments

OpenStudy (chrisplusian):

A trigonometric fourier series is represented: x(t)=2+3cos(2t)+4sin(2t)+2sin(3t+30)-cos(7t-150) The Compact form is : 3cos(2t)+4sin(2t)=5cos(2t-53.13) *AS a side note the phase shifts here are in degrees

OpenStudy (chrisplusian):

I just want to understand how they manipulated one into the other

OpenStudy (thomas5267):

\[ A\cos(a+b)=A\cos(a)\cos(b)-A\sin(a)\sin(b)\\ A\cos(2t+b)=A\cos(2t)\cos(b)-A\sin(2t)\sin(b)\\ A\cos(b)=3\\ A\sin(b)=4\\ \tan(b)=\frac{4}{3}\\ A^2\cos^2(b)+A^2\sin^2(b)=3^2+4^2=25\\ A=5 \]